OpenStudy (anonymous):

in triangle PQR, the measure of angle p is 60, the measure of angle q is 30, and the measure of angle r in 90. which of the following statements of pqr are true. check all A. QR=(root3/2)*PQ B. PR=(root3/2)*PQ C. QR=root3*PR

OpenStudy (anonymous):

D. PQ=root3*PR E. QR=2*PR F. PQ=2*PR